Suppose you loaned your grandmother $1000 today. You let her have the money for 1% interest for two years. Unfortunately, inflation is 2%. How much money does grandma give you in 2 years? What is the current purchasing power of your money? Show your calculations.
Answer #1

Grandmother will give amount=P(1+r)t=1000(1.01)2=1020.1

and inflation is 2% in 2 yearr. Thus net increase in price=4%

new price=1.04old price

Thus purchasing power=1020.1/1.04=980.86